Suppose a consumer’s utility function is given a U = 100X0.25Y0.75.The prices of the two commodities and are Birr 2 and Birr 5 per unit respectively. If the consumer’s income isBirr 280, how many units of each commodity should the consumer buy to maximize his/her utility? 


1
Expert's answer
2021-08-03T11:50:26-0400

Solution:

Utility maximizing condition:  "\frac{MUx}{MUy} = \frac{Px}{Py}"

MRS = "\frac{MUx}{MUy} = \frac{Px}{Py}"

First, derive MRS:

MRS = "\frac{MUx}{MUy} = \frac{Px}{Py}"


MUx =  "\frac{\partial U} {\partial X}" = 25Y0.75


MUy = "\frac{\partial U} {\partial Y} =" 75X0.25


MRS = "\frac{MUx}{MUy}" = "\frac{25Y^{0.75} }{75X^{0.25} }"


Set MRS equal to Px/Py to derive the utility maximizing bundle:

Px = 2

Py = 5

"\frac{25Y^{0.75} }{75X^{0.25} }" = "\frac{2}{5}"


X = "\frac{625Y^{3} }{1296 }"


Plug X into the budget constraint to derive Y:

Budget constraint: M = PxX + PyY

280 = 2X + 5Y

280 = 2(625Y3/1296) + 5Y

Y = 6.4

Plug this into X equation:

X = 625Y3/1296 = 625(6.43)/1296 = 163840/1296 = 126.4

X = 126.4

Utility maximizing bundle (Ux,y) = (126.4, 6.4)